# Generative Deep Learning Framework for Inverse Design of Fuels

### v2

@@ Abstract
Fuel design , Generative Deep Learning , Variational Autoencoder , Research Octane Number[label1] organization=Transportation and Power Systems Division, Argonne National Laboratory, addressline=9700 S Cass Avenue, city=Lemont, postcode=IL 60439, country=USA
[label2] organization=Digital R&D Division, Research and Development Center, Saudi Aramco, city=Dhahran, country=Saudi Arabia
[label3] organization=Transport Technologies Division, Research and Development Center, Saudi Aramco, city=Dhahran, country=Saudi Arabia
[label4] organization=Aramco Americas: Aramco Research Center–Detroit, addressline=Novi, MI, country=USA
@@ 1 Introduction
Generative deep learning models have significantly transformed the exploration and design of new chemical structures, particularly in the pharmaceutical industry. Models such as Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), and Reinforcement Learning (RL) frameworks have demonstrated a remarkable ability to generate novel, drug-like molecules with desired properties [1, 2, 3, 4]. Inspired by these successes, researchers have begun applying generative models to fuel molecular design. Liu et al. [5] introduced a generative fuel design framework effective in low-data scenarios by employing multi-objective imitation learning. This suggests that generative models can be tailored to address fuel design challenges where data on candidate molecule properties is scarce. Furthermore, Rittig et al. [6] presented a graph-based generative modeling strategy integrated with optimization algorithms to design high-octane fuels. Their graph machine-learning framework successfully identified known high-octane compounds. It proposed new fuel candidates, highlighting the growing potential of generative AI in discovering fuel molecules that meet specific performance requirements. Kuzhagaliyeva et al. [7] expanded this concept to fuel mixtures, utilizing AI to design fuel blends with optimal performance, further demonstrating the versatility of AI-driven design in the combustion field. Recently, Yalamanchi et al. [8] further enhanced this model by incorporating an uncertainty quantification (UQ) framework.
@@
Simultaneously, significant progress has been made in machine learning-based Quantitative Structure-Property Relationship (QSPR) models for predicting fuel properties. Traditional QSPR methods often employ group contribution techniques or physicochemical descriptors to estimate metrics such as octane numbers. In recent years, data-driven models using molecular fingerprints and graph representations have achieved improved accuracy in predicting octane numbers and sensitivities. Vom Lehn et al. [9] showed that tailoring feature sets for specific targets—such as Research Octane Number (RON) and Motor Octane Number (MON)—and selecting appropriate machine learning algorithms, including neural networks, enhanced predictive performance for gasoline compounds. SubLaban et al. [10] developed artificial neural network models for RON and octane sensitivity, providing a QSPR approach for fuel design. Their models underscored the effectiveness of nonlinear machine learning in capturing complex structure-property relationships beyond the capabilities of linear correlations. Additionally, Mohan and Chang [11] introduced an automated machine learning framework called ChemSL, which evaluates multiple algorithms and hyperparameters to construct robust models for property prediction. Such automated and optimized approaches are particularly relevant when integrated with generative models, ensuring the predictive component maintains high accuracy.
@@
These studies suggest that generative models, such as VAEs, can learn meaningful molecular representations that are useful for proposing new fuel molecules, and machine learning-based QSPR models can accurately predict key fuel properties when properly tuned. However, while advanced machine learning methods can enhance prediction accuracy, many QSPR models are trained on relatively limited datasets with hand-crafted features, which may not generalize across broad chemical spaces. This limitation introduces the risk of overlooking promising novel compounds outside the model’s experience. There is a growing interest in integrating generative and predictive modeling for fuel design to overcome these challenges. Researchers can efficiently navigate the chemical space to identify molecules with desired fuel properties by coupling a generative model with a property predictor. By training generative models on comprehensive chemical databases that cover a vast chemical space and applying property predictions to the latent representations or generated candidates, researchers can efficiently conduct targeted searches for molecules with desired properties. Liu et al. [12] developed a VAE architecture that jointly optimized the compression objective and property prediction within a single integrated framework. This demonstrates the model’s ability to capture essential chemical features relevant to predicting heating values and ignition metrics, even when property data were limited. These insights inform the approach adopted in this work, which builds on this concept by separating the generative and predictive components into distinct models. This modular design enhances flexibility, allowing independent training and optimization of the VAE-based generative model and the data-driven property predictor, ultimately improving the robustness and performance of each component.
@@
The authors’ previous work [13] developed a VAE-based framework that learns a continuous latent representation of molecular structures suitable for generation and prediction. The VAE was trained on a downsampled GDB-13 chemical database using one-hot-encoded Simplified Molecular Input Line Entry System (SMILES) strings as input. Strategies to balance reconstruction fidelity and latent space regularization through the weightage of the regularization term (β) were systematically investigated. It was found that gradually increasing β to a moderate value (0.25) provided the best trade-off between reconstruction accuracy and latent space consistency. Higher β values risked posterior collapse—where the decoder ignores latent features—while lower β values led to poor regularization. Advanced techniques, such as cyclical β schedules [14], hierarchical priors [15], and penalizing latent variable correlations [16], were also explored. The results indicated that although a hierarchical latent prior offered some improvements in reconstruction accuracy, its benefits were modest, and substantial penalization of latent total correlation degraded performance. Employing a standard VAE with an annealed β schedule up to 0.25 captured most of the improvements.
@@
Building upon these findings, the current study focuses on a streamlined yet enhanced VAE-based model for inverse fuel design. To this end, a co-optimized VAE (Co-VAE) approach is introduced that directly incorporates fuel property information into the generative model, aiming to steer the latent space toward meaningful features for property prediction. This Co-VAE is trained using the previously optimized β-annealing schedule, which increases β from zero to 0.25 over 75 epochs without introducing additional complexities, such as hierarchical priors or total-correlation regularization. Concurrently, rigorous hyperparameter optimization (HPO) of the Co-VAE architecture and training process is performed using Bayesian optimization to ensure the model’s capacity and learning dynamics are tuned to the data. Additionally, a QSPR modeling step is integrated by mapping the learned latent representations to the target fuel property (Research Octane Number (RON)) constructing separate regression models for fine-tuning. Multiple machine learning algorithms are evaluated for this task, applying systematic hyperparameter searches to identify the optimal predictor. Subsequently, a latent space screening approach is used to pinpoint molecular candidates exhibiting promising fuel property. Combining the optimized predictive model with a robust global search strategy and thorough chemical validity checks, a diverse set of high-RON molecules are obtained for further investigation.
@@ 2.1 Data Curation and Processing
The datasets used in this study are fundamental to developing the Co-VAE model and accurately predicting RON. Building upon the authors’ previous work [13], a refined subset of the GDB-13 database [17] was employed, which is an extensive collection of small organic molecules containing up to 13 heavy atoms encompassing over 970 million unique compounds. The subset was downselected by focusing on molecules relevant to fuel applications, tailoring the dataset to the specific objective of fuel design, and managing computational feasibility. Specifically, compounds composed solely of carbon (C), hydrogen (H), and oxygen (O) atoms were selected, and the molecular size was limited to those containing up to 10 heavy atoms. This initial refinement yielded a dataset of 418,274 unique species, offering a manageable yet diverse representation of potential fuel molecules. In the present work, this dataset was further pruned by removing species containing more than two cyclic rings. Highly polycyclic compounds introduce significant structural complexity and are less pertinent to practical fuel formulations. This additional filtering ensures that the dataset remains focused on molecules that are both computationally tractable and relevant to fuel chemistry.
@@
Experimental RON data was further integrated to enhance the dataset’s applicability to property prediction tasks. A comprehensive RON dataset was curated from various literature sources [7, 18, 19], all of which compile ASTM-compliant RON values from direct experimental measurements, selecting the most recent and reliable measurements for single compounds. This careful curation, applying the same down-sampling criteria as those used for the GDB-13 dataset, resulted in a dataset comprising 332 species along with associated RON values. Some species in the RON dataset were not present in the downselected GDB-13 subset. These missing species were incorporated into the dataset to ensure comprehensive coverage and maximize the model’s ability to learn structure-property relationships. With this incorporation, the final downsized dataset comprises 357,907 species.
@@
Molecular structures were represented using the SMILES notation, which encodes complex molecular graphs into linear strings suitable for computational processing. These SMILES strings were converted into one-hot encoded representations to serve as inputs to the Co-VAE model. The one-hot encoding process involved mapping each character in the SMILES string to a unique position in a binary vector, resulting in a fixed-length, two-dimensional array. All SMILES strings were padded to a maximum length of 23 characters to maintain consistency, and the character set comprised eight unique symbols representing atoms and molecular substructures [13]. This encoding ensures that the input data is uniform in size and compatible with the Long Short-Term Memory (LSTM) networks employed in the Co-VAE architecture (discussed in the next subsection). Using one-hot encoding for SMILES strings facilitates the direct reconstruction of molecular structures from the latent space embeddings generated by the Co-VAE model. This approach offers advantages over traditional molecular descriptor-based methods by enabling an interpretable transformation between molecular structures and their encoded representations without requiring similarity search algorithms.
@@ 2.2 Model Architecture
The Co-VAE model developed in this study builds upon the authors’ previous VAE framework, designed for molecular generation and property prediction. The original VAE architecture consists of a two-layer Long Short-Term Memory (LSTM) network [20], serving as the encoder, which is chosen for its ability to capture sequential dependencies in one-hot encoded SMILES strings that represent molecular structures. The encoder processes the input sequences and funnels the information through two fully connected (FC) layers into the latent space, producing means and log-variances that define the distribution from which latent variables are sampled. The decoder mirrors this structure in reverse, expanding the latent variables back to their original dimensionality through fully connected (FC) layers and reconstructing the input sequences using an LSTM network over the sequence length. For a more detailed description of this architecture, the readers are referred to authors’ previous work [13].
@@
where is the weight balancing the contribution of the KLD term, and represents the mean absolute error (MAE) associated with RON prediction.
@@ 2.3 Hyperparameter Optimization
Optimizing the hyperparameters of the Co-VAE model is crucial for enhancing its performance in both molecular generation and property prediction tasks. This study systematically explored various hyperparameters that influence the neural network’s capacity, the complexity of the latent space, and the efficiency of the training process. The hyperparameters varied and included the number of layers in the LSTM encoder and decoder, the hidden size of the LSTM layers, the sizes of the fully connected and conditioning layers, the dimensionality of the latent space, and the batch size. The specific ranges within which these hyperparameters were varied are provided in Table 1. The learning rate remained fixed at 1×10-3 and the number of training epochs was set to 300. Preliminary experiments indicated that the model consistently converged within this epoch range, justifying the fixed number of epochs and learning rate. The architecture’s capacity and latent space properties was optimized to fine-tune the model, enabling it to capture the essential features necessary for accurate molecular reconstruction and RON prediction.
@@
The down-selected GDB-13 dataset was split into 95 for training, 2.5 for validation, and 2.5 for testing. For the RON dataset, 10 species were selected for the validation set and 10 for the test set. The BCE and KLD terms in the loss function are computed for the entire GDB-derived dataset, regardless of whether or not RON measurements are available. However, only the molecules with experimentally measured RON values contribute to . Specifically, when a molecule in a training batch has a known RON value, its RON prediction error is included alongside the BCE and KLD loss terms; for other molecules, only the BCE and KLD loss terms are non-zero. To efficiently navigate the hyperparameter space, Bayesian optimization using the bayesopt package [21] was employed. This methodology is well-suited for optimizing complex functions where evaluation is costly, as it intelligently selects hyperparameter configurations based on previous results to improve performance iteratively. The optimization process was initiated with 16 random evaluations to broadly explore diverse regions of the hyperparameter space. Following this initial exploration, sequential optimization steps were conducted, selecting 8 points per batch and performing 40 runs. This approach enabled us to progressively refine the hyperparameters based on the observed model performance. The optimization metric was defined as a composite score based on the validation set, combining reconstruction accuracy, validity of generated molecules, and loss associated with predicting RON. Reconstruction accuracy measures the model’s ability to recreate input SMILES strings from the latent representations accurately. Validity assesses the proportion of generated molecules from the sampled normal distribution for latent space that is chemically plausible and adheres to established chemical valency rules, ensuring that outputs are meaningful within the domain of chemistry. Including the RON prediction loss in the metric encourages the latent space to be informative for property prediction, aligning the generative model with the target predictive task.
@@
The results of the HPO, derived from the validation set, are depicted in Figure 3, which illustrates the progression of the optimization metric over iterations and demonstrates convergence toward the optimal configuration. The composite score is computed as the sum of reconstruction accuracy, validity and five times the inverse of RON MAE to ensure comparable magnitudes. Upon completing the optimization runs, the hyperparameter configuration that yielded the best composite score on the validation set was selected as the optimal model, corresponding to iteration 16 in Figure 3. The specific values for each hyperparameter for this optimal configuration and the corresponding results on the test set are summarized in Table 2.

### v3

@@ Abstract
[…]
@@ 1 Introduction
Generative deep learning models have significantly transformed the exploration and design of new chemical structures, particularly in the pharmaceutical industry. Models such as Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), and Reinforcement Learning (RL) frameworks have demonstrated a remarkable ability to generate novel, drug-like molecules with desired properties 1, 2, 3, 4. Inspired by these successes, researchers have begun applying generative models to fuel molecular design. Liu et al. 5 introduced a generative fuel design framework effective in low-data scenarios by employing multi-objective imitation learning. This suggests that generative models can be tailored to address fuel design challenges where data on candidate molecule properties is scarce. Furthermore, Rittig et al. 6 presented a graph-based generative modeling strategy integrated with optimization algorithms to design high-octane fuels. Their graph machine-learning framework successfully identified known high-octane compounds. It proposed new fuel candidates, highlighting the growing potential of generative AI in discovering fuel molecules that meet specific performance requirements. Furthermore, recent work 7 has employed LSTM-based generative model fine-tuned with hill-climb optimization to generate novel fuel molecules specifically optimized for high-knock resistance, demonstrating how generative approaches can be tailored to produce molecules with targeted fuel properties. Kuzhagaliyeva et al. 8 expanded this concept to fuel mixtures, utilizing AI to design fuel blends with optimal performance, further demonstrating the versatility of AI-driven design in the combustion field. Recently, Yalamanchi et al. 9 further enhanced this model by incorporating an uncertainty quantification (UQ) framework.
@@
Simultaneously, significant progress has been made in machine learning-based Quantitative Structure-Property Relationship (QSPR) models for predicting fuel properties. Traditional QSPR methods often employ group contribution techniques or physicochemical descriptors to estimate metrics such as octane numbers. In recent years, data-driven models using molecular fingerprints and graph representations have achieved improved accuracy in predicting octane numbers and sensitivities. Vom Lehn et al. 10 showed that tailoring feature sets for specific targets—such as Research Octane Number (RON) and Motor Octane Number (MON)—and selecting appropriate machine learning algorithms, including neural networks, enhanced predictive performance for gasoline compounds. SubLaban et al. 11 developed artificial neural network models for RON and octane sensitivity, providing a QSPR approach for fuel design. Their models underscored the effectiveness of nonlinear machine learning in capturing complex structure-property relationships beyond the capabilities of linear correlations. Additionally, Mohan and Chang 12 introduced an automated machine learning framework called ChemSL, which evaluates multiple algorithms and hyperparameters to construct robust models for property prediction. Such automated and optimized approaches are particularly relevant when integrated with generative models, ensuring the predictive component maintains high accuracy. Chen et al. 13 demonstrated a comprehensive approach of screening millions of fuel candidates using QSPR models to identify molecules tailor-made for specific applications, highlighting how these predictive models can be used to evaluate the practical application potential of vast molecular libraries. However, these QSPR-based approaches differ fundamentally from generative methods by focusing on property prediction and screening rather than molecular generation. While QSPR models excel at evaluating properties of known or proposed structures, they do not inherently generate novel molecular candidates, highlighting a key methodological distinction from generative approaches that can explore uncharted regions of chemical space.
@@
These studies suggest that generative models, such as VAEs, can learn meaningful molecular representations that are useful for proposing new fuel molecules, and machine learning-based QSPR models can accurately predict key fuel properties when properly tuned. However, while advanced machine learning methods can enhance prediction accuracy, many QSPR models are trained on relatively limited datasets with hand-crafted features, which may not generalize across broad chemical spaces. This limitation introduces the risk of overlooking promising novel compounds outside the model’s experience. There is a growing interest in integrating generative and predictive modeling for fuel design to overcome these challenges. Researchers can efficiently navigate the chemical space to identify molecules with desired fuel properties by coupling a generative model with a property predictor. By training generative models on comprehensive chemical databases that cover a vast chemical space and applying property predictions to the latent representations or generated candidates, researchers can efficiently conduct targeted searches for molecules with desired properties. Liu et al. 14 developed a VAE architecture that jointly optimized the compression objective and property prediction within a single integrated framework. This demonstrates the model’s ability to capture essential chemical features relevant to predicting heating values and ignition metrics, even when property data were limited. These insights inform the approach adopted in this work, which builds on this concept by separating the generative and predictive components into distinct models. This modular design enhances flexibility, allowing independent training and optimization of the VAE-based generative model and the data-driven property predictor, ultimately improving the robustness and performance of each component.
@@
The authors’ previous work 15 developed a VAE-based framework that learns a continuous latent representation of molecular structures suitable for generation and prediction. The VAE was trained on a downsampled GDB-13 chemical database using one-hot-encoded Simplified Molecular Input Line Entry System (SMILES) strings as input. Strategies to balance reconstruction fidelity and latent space regularization through the weightage of the regularization term (β) were systematically investigated. It was found that gradually increasing β to a moderate value (0.25) provided the best trade-off between reconstruction accuracy and latent space consistency. Higher β values risked posterior collapse—where the decoder ignores latent features—while lower β values led to poor regularization. Advanced techniques, such as cyclical β schedules 16, hierarchical priors 17, and penalizing latent variable correlations 18, were also explored. The results indicated that although a hierarchical latent prior offered some improvements in reconstruction accuracy, its benefits were modest, and substantial penalization of latent total correlation degraded performance. Employing a standard VAE with an annealed β schedule up to 0.25 captured most of the improvements.
@@
Building upon these findings, the current study focuses on a streamlined yet enhanced VAE-based model for inverse fuel design. To this end, a co-optimized VAE (Co-VAE) approach is introduced that directly incorporates fuel property information into the generative model, aiming to steer the latent space toward meaningful features for property prediction. This Co-VAE is trained using the previously optimized β-annealing schedule, which increases β from zero to 0.25 over 75 epochs without introducing additional complexities, such as hierarchical priors or total-correlation regularization. Concurrently, rigorous hyperparameter optimization (HPO) of the Co-VAE architecture and training process is performed using Bayesian optimization to ensure that the model’s capacity and learning dynamics are tuned to the data. Additionally, a QSPR modeling step is integrated by mapping the learned latent representations to the target fuel property (Research Octane Number - RON) constructing separate regression models for fine-tuning. Multiple machine learning algorithms are evaluated for this task, applying systematic hyperparameter searches to identify the optimal predictor. Subsequently, a latent space screening approach is used to pinpoint molecular candidates exhibiting promising fuel property. Combining the optimized predictive model with a robust global search strategy and thorough chemical validity checks, a diverse set of high-RON molecules are obtained for further investigation.
@@ 2.1 Data Curation and Processing
The datasets used in this study are fundamental to developing the Co-VAE model and accurately predicting RON. Building upon the authors’ previous work 15, a refined subset of the GDB-13 database 19 was employed, which is an extensive collection of small organic molecules containing up to 13 heavy atoms encompassing over 970 million unique compounds. The subset was downselected by focusing on molecules relevant to fuel applications, tailoring the dataset to the specific objective of fuel design, and managing computational feasibility. Specifically, compounds composed solely of carbon (C), hydrogen (H), and oxygen (O) atoms were selected, and the molecular size was limited to those containing up to 10 heavy atoms. This initial refinement yielded a dataset of 418,274 unique species, offering a manageable yet diverse representation of potential fuel molecules. In the present work, this dataset was further pruned by removing species containing more than two cyclic rings. Highly polycyclic compounds introduce significant structural complexity and are less pertinent to practical fuel formulations. This additional filtering ensures that the dataset remains focused on molecules that are both computationally tractable and relevant to fuel chemistry.
@@
Experimental RON data was further integrated to enhance the dataset’s applicability to property prediction tasks. A comprehensive RON dataset was curated from literature sources 8, 20, which compile ASTM-compliant RON values, most of which originate from previously reported experimental measurements 21. This careful curation, applying the same down-sampling criteria as those used for the GDB-13 dataset, resulted in a dataset comprising 332 species along with associated RON values. Although the downselected GDB dataset is large, it is still not exhaustive of all fuel-like species and therefore missed out some species from the RON dataset. These missing species were incorporated into the dataset to ensure comprehensive coverage and maximize the model’s ability to learn structure-property relationships. With this incorporation, the final downsized dataset comprises 357,907 species.
@@
Molecular structures were represented using the SMILES notation, which encodes complex molecular graphs into linear strings suitable for computational processing. These SMILES strings were converted into one-hot encoded representations to serve as inputs to the Co-VAE model. The one-hot encoding process involved mapping each character in the SMILES string to a unique position in a binary vector, resulting in a fixed-length, two-dimensional array. All SMILES strings were padded to a maximum length of 23 characters to maintain consistency, and the character set comprised eight unique symbols representing atoms and molecular substructures 15. This encoding ensures that the input data is uniform in size and compatible with the Long Short-Term Memory (LSTM) networks employed in the Co-VAE architecture (discussed in the next subsection). Using one-hot encoding for SMILES strings facilitates the direct reconstruction of molecular structures from the latent space embeddings generated by the Co-VAE model. This approach offers advantages over traditional molecular descriptor-based methods by enabling an interpretable transformation between molecular structures and their encoded representations without requiring similarity search algorithms.
@@ 2.2 Model Architecture
The Co-VAE model developed in this study builds upon the authors’ previous VAE framework, designed for molecular generation and property prediction. The original VAE architecture consists of a two-layer Long Short-Term Memory (LSTM) network 22, serving as the encoder, which is chosen for its ability to capture sequential dependencies in one-hot encoded SMILES strings that represent molecular structures. The encoder processes the input sequences and funnels the information through two fully connected (FC) layers into the latent space, producing means and log-variances that define the distribution from which latent variables are sampled. The decoder mirrors this structure in reverse, expanding the latent variables back to their original dimensionality through fully connected (FC) layers and reconstructing the input sequences using an LSTM network over the sequence length. For a more detailed description of this architecture, the readers are referred to authors’ previous work 15.
@@
where is the weight balancing the contribution of the KLD term. BCE quantifies the reconstruction error by measuring the cross-entropy between predicted and target binary values in the SMILES strings, while KLD regularizes the latent space by penalizing deviations from a standard normal distribution. represents the mean absolute error (MAE) associated with RON prediction.
@@ 2.3 Hyperparameter Optimization
Optimizing the hyperparameters of the Co-VAE model is crucial for enhancing its performance in both molecular generation and property prediction tasks. This study systematically explored various hyperparameters that influence the neural network’s capacity, the complexity of the latent space, and the efficiency of the training process. The hyperparameters varied and included the number of layers in the LSTM encoder and decoder, the hidden size of the LSTM layers, the sizes of the fully connected and conditioning layers, the dimensionality of the latent space, and the batch size. The specific ranges within which these hyperparameters were varied are provided in Table 1. The learning rate remained fixed at 1× and the number of training epochs was set to 300. Preliminary experiments indicated that the model consistently converged within this epoch range, justifying the fixed number of epochs and learning rate. The architecture’s capacity and latent space properties were optimized to fine-tune the model, enabling it to capture the essential features necessary for accurate molecular reconstruction and RON prediction.
@@
The downselected GDB-13 dataset was split into 95 for training, 2.5 for validation, and 2.5 for testing. For the RON dataset, 10 species were selected for the validation set and 10 for the test set. To ensure statistical diversity, stratified sampling was employed based on molecular size and functional groups for the GDB-13 dataset, and based on molecular size for the RON dataset. The BCE and KLD terms in the loss function are computed for the entire GDB-derived dataset, regardless of whether or not RON measurements are available. However, only the molecules with experimentally measured RON values contribute to . Specifically, when a molecule in a training batch has a known RON value, its RON prediction error is included alongside the BCE and KLD loss terms; for other molecules, only the BCE and KLD loss terms are non-zero. To efficiently navigate the hyperparameter space, Bayesian optimization using the bayesopt package 23 was employed. This methodology is well-suited for optimizing complex functions where evaluation is costly, as it intelligently selects hyperparameter configurations based on previous results to improve performance iteratively. The optimization process was initiated with 16 random evaluations to broadly explore diverse regions of the hyperparameter space. Following this initial exploration, sequential optimization steps were conducted, selecting 8 points per batch and performing 40 runs. This approach enabled us to progressively refine the hyperparameters based on the observed model performance. The optimization metric was defined as a composite score based on the validation set, combining reconstruction accuracy, validity of generated molecules, and loss associated with predicting RON. Reconstruction accuracy measures the model’s ability to recreate input SMILES strings from the latent representations accurately. Validity assesses the proportion of generated molecules from the sampled normal distribution for latent space that is chemically plausible and adheres to established chemical valency rules, ensuring that outputs are meaningful within the domain of chemistry. Including the RON prediction loss in the metric encourages the latent space to be informative for property prediction, aligning the generative model with the target predictive task.
@@
The results of the HPO, derived from the validation set, are depicted in Figure 3, which illustrates the optimization metrics over iterations. Different configurations yielded varying performance levels across the optimization runs. The composite score is computed as the sum of reconstruction accuracy, validity and five times the inverse of RON MAE to ensure comparable magnitudes. Upon completing the optimization runs, the hyperparameter configuration that yielded the best composite score on the validation set was selected as the optimal model, corresponding to iteration 16 in Figure 3. The specific values for each hyperparameter for this optimal configuration and the corresponding results on the test set are summarized in Table 2. The relatively high MAE of RON prediction from the Co-VAE model can be attributed to the significant difference in dataset sizes, with the RON dataset being orders of magnitude smaller than the downselected GDB-13 dataset. This limitation necessitates the development of a separate regression model, which will be discussed in the following section.